Suppose that in addition to $5.55 million of taxable income, Texas Taco Inc., received $600,000 of interest on state-issued bonds and 260,000 of dividends of common stock it owns in Arizona Taco, Inc., (Use corporate tax rate of 21 percent for your calculations.)

1. What is Texas Taco's income tax liability? (Answers should be in dollars not millions)

2. What are Texas Taco's average and marginal tax rates on taxable income? (Round answers 2 decimal places)
Answer #1
1] Taxable income $    55,50,000
Add: Taxable dividends = 260000*30% = $          78,000
Net taxable income $    56,28,000
Income tax liability = 5628000*21% =- $    11,81,880
2] Average rate of tax 21.00%
Marginal rate of tax 21.00%